The Nelson Company has $1,400,000 in current assets and $500,000 in current liabilities. Its initial inventory level is $350,000, and it will raise funds as additional notes payable and use them to increase inventory.
1. How much can Nelson's short-term debt (notes payable) increase without pushing its current ratio below 1.4? Round your answer to the nearest cent. $________
2. What will be the firm's quick ratio after Nelson has raised the maximum amount of short-term funds? Round your answer to two decimal places. ________
Answer #1

Current Ratio = Initial Current Assets /(Current Iiabilities + Short Term Debt) = 1.4
1,400,000/(500,000+Short Term Debt) = 1.4
1,400,000 = 1.4 * Short term debt + 700,000
Short Term Debt = (1400000-700000)/1.4 = 500,000

2. Quick Ratio after short term fund is raised = (Initial Current Assets - Inventories)/(Current Iiabilities + Short Term Debt)
=(1400000-350000)/(500000+500000) = 1.05